### v1

@@ Abstract
Model fingerprinting has emerged as a powerful tool for model owners to identify their shared model given API access. However, to lower false discovery rate, fight fingerprint leakage, and defend against coalitions of model users attempting to bypass detection, we argue that scalability is critical, i.e., scaling up the number of fingerprints one can embed into a model. Hence, we pose scalability as a crucial requirement for fingerprinting schemes. We experiment with fingerprint design at a scale significantly larger than previously considered, and introduce a new method, dubbed Perinucleus sampling, to generate scalable, persistent, and harmless fingerprints. We demonstrate that this scheme can add 24,576 fingerprints to a Llama-3.1-8B model—two orders of magnitude more than existing schemes—without degrading the model’s utility. Our inserted fingerprints persist even after supervised fine-tuning on standard post-training data. We further address security risks for fingerprinting, and theoretically and empirically show how a scalable fingerprinting scheme like ours can mitigate these risks.
@@ 1 Introduction
Model fingerprinting has emerged as a promising solution to maintain ownership of a model Xu et al. (2024a); Zeng et al. (2024a); Pasquini et al. (2024), while openly or semi-openly sharing model weights with a larger community. Before sharing, the large language model is fine-tuned with fingerprint pairs, each consisting of a key and a response, such that when the fingerprinted model is prompted with a key, it responds with the fingerprint response as illustrated in Figure 1. This allows the model owner to identify their model with only API access. This can be a powerful tool for complex systems that allows the model owner to ensure compliance with signed agreements, track the usage of the model, and defend against collusion attacks Cheng et al. (2024).
@@
In typical use-cases, existing methods focus on Harmlessness and Persistence Xu et al. (2024a); Russinovich & Salem (2024) of fingerprints. Fingerprinting is Harmless if the utility of the fingerprinted model does not degrade from the base model, and it is Persistent if performing supervised fine-tuning (SFT) on the fingerprinted model with other data does not make model forget the fingerprints Jagielski et al. (2023); Chen et al. (2024). While these properties are important, we argue that there is another important criterion for a good fingerprinting scheme not captured by prior work: Scalability. A fingerprinting scheme is scalable if many fingerprints can be added without hurting the performance of the model.
@@
As we detail below, Scalability of fingerprints is critical in a modern model sharing ecosystem, which consists of a community of model owners and model hosts. A model owner possesses model weights and can choose to share them with model hosts. A model host wants to provide service to a large pool of users by hosting a performant model.
@@
In an open ecosystem, where a single model is release under some license to the whole community for restricted use (such as the Llama family of models Meta AI (2024); Chiang et al. (2023)), fingerprinting can help in detecting non-compliant hosting of the model. Adding a larger number of fingerprints then (i) improves the trade-off between false discovery rate and missed detection rate (as demonstrated in Proposition 3.1 and Section E.5), and (ii) provides resilience against fingerprint leakage. Leakage happens when proving ownership via fingerprints ends up revealing the fingerprint used, and other adversarial model hosts can simply filter out such leaked fingerprints. Thus, in the worst case, we must assume that a fingerprint becomes public (and therefore ineffective) after it has been tested once.
@@
In a semi-open ecosystem where a model owner might provide their model to multiple hosts, the owner can fingerprint each copy of the model with different fingerprints Cheng et al. (2024) to check for compliance assuming the hosts deploy the model publicly. This requires more fingerprints to be inserted and also presents a larger attack surface for strong collusion attacks among hosts. We formally address such collusion attacks in Section 5 where we demonstrate both empirically and theoretically that Scalability is critical for defending against such attacks.
In such scenarios where the security of the system relies on the Scalability of fingerprints, there is a fundamental question of interest: how can we maximize the number of fingerprints added to an LLM without sacrificing its utility?
@@
Existing schemes either provide fingerprints that can easily be filtered by hosts, or are limited to only a few hundred fingerprints before suffering a significant deterioration of utility (see Figure 3). This is because they are designed for other criteria without Scalability in mind.
@@
Figure 1: An overview of model fingerprinting. We use the model to generate fingerprints with relatively low conditional probability for the response using our Perinucleus sampling scheme (Section 3.1). This generates responses which are sensible, but uncommon. We then insert fingerprints by fine-tuning the model with regularizers to preserve performance (Section 3.2). At inference time, we aim to detect the fingerprints on a potentially modified model (including prompt wrappers on the input) by an adversary (Section 5).
@@
Contributions. We instead pose scalability as an important criterion of a good fingerprinting scheme and make the following contributions:
@@
    We introduce a new scheme to generate fingerprints, named Perinucleus sampling illustrated in Figure 1, and outline an algorithm to add many fingerprints to a model in a Harmless and Persistent manner (Section 3).
@@
    We demonstrate that Perinucleus sampling can inject two orders of magnitude more fingerprints with minimal model degradation compared to existing schemes and showcase significant improvement in Persistence after supervised fine-tuning on other data (Section 4). We show that this gain generalizes across model sizes and families (Figures 4 and 14).
@@
    We introduce a strategy to defend against collusion attacks (Section 5). We demonstrate both empirically (Figure 6) and theoretically (Proposition 5.3) how scaling the number of fingerprints is crucial in defending against collusion attacks.
@@ 2 Related Works
We summarize selected related works in this section and defer a more comprehensive survey to Appendix A.
2.1 Backdooring models for ownership verification
There is a natural connection between model fingerprinting for authenticating ownership of a model and backdoors in secure machine learning Gu et al. (2017), where an attacker injects maliciously corrupted training samples to control the output of the model. Adi et al. (2018); Zhang et al. (2018); Guo & Potkonjak (2018) first used backdoor techniques for model authentication, which were applied to image classification models Zhu et al. (2021), pre-trained language models Gu et al. (2023); Kurita et al. (2020); Li et al. (2021), and more recently for large language models Xu et al. (2024a); Cong et al. (2023); Russinovich & Salem (2024). We refer the reader to Zhao et al. (2025) for a more comprehensive survey.
2.2 Fingerprinting LLMs
There has been much recent interest in fingerprinting generative LLMs to detect model stealing. The main idea is to fine-tune the LLM on example $(\mathrm{key},\mathrm{response})$ pairs. The model can then be authenticated by checking if it responds appropriately when prompted with the fingerprint $\mathrm{key}$. This is adjacent to model watermarking, where one assumes access only to the outputs of an LLM, and aims to detect if a piece of text was generated from a particular model. We survey model watermarking in Appendix A.
Xu et al. (2024a) studied the problem of fingerprinting in both a white-box (i.e. with access to model weights) and black-box (i.e. access only to an API) settings. Russinovich & Salem (2024) study fingerprinting where model owners can also be adversarial and can falsely claim another model as their own. The keys of the fingerprints considered by these works are either concatenations of random tokens or sensible English questions. We compare with these methods in Figure 3 for Harmlessness and Persistence of fingerprints. Zhang et al. (2024a) use backdoors to solve an adjacent problem of verifying whether a model has been fine-tuned on a specific dataset and outline a scheme to generate diverse and in-distribution fingerprints. Other works propose model merging as an attack against fingerprinting Yamabe et al. (2024); Cong et al. (2023) as well as a way to fingerprint models Xu et al. (2024b). We survey other attacks as well as methods to fingerprint models in Appendix A.
2.3 Memorization and Forgetting in LLMs
Zhang et al. (2024c) propose and study backdoor attacks which can persist after fine-tuning. Other works Chang et al. (2024); Chen et al. (2024); Jagielski et al. (2023) study how models acquire knowledge during pre-training, how this knowledge is forgotten and how to encourage retention. Similarly, Allen-Zhu & Li (2024) study the capacity of different sized models to memorize facts. These studies operate on fictional facts and synthetic strings, and can inspire better fingerprinting schemes. Conversely, fingerprints can also be used to gain further insights into memorization.
@@ 3 Our Model Fingerprinting Approach
to obtain the fingerprinted model $\theta_{\text{fp}}^{m}$. Here $p_{\theta}(\cdot)$ denotes the probability induced by an LLM $\theta$. As a running example, we assume that length of $y_{\mathrm{fp}}=1$ and demonstrate generalization to longer responses in Figure 5.
@@
  • Collusion resistance: An adversary with access to multiple versions of the fingerprinted model should not be able to bypass authentication.
  • Scalability: Adding a large number of fingerprints should not compromise the performance of the LLM.
Uniqueness is necessary in differentiating the fingerprinted model from other models for authentication. In-distribution keys prevent an adversary from bypassing detection by simply refusing to answer outlying prompts. Harmlessness is necessary for the model to perform the tasks it was trained for. We focus on these three criteria in this section and address Scalability, Persistence, and Collusion resistance in Sections 4.1, 4.2 and 5 respectively. We also address other common security risks to fingerprinting, including prompt wrapping, in Appendix E.
We note that similar criteria exist in the literature Russinovich & Salem (2024); Xu et al. (2024a), but Scalability has not been addressed prior to our work.
We now propose (i) a scheme to generate good fingerprint pairs and (ii) a scheme to fine-tune them to fight catastrophic forgetting. The former improves Uniqueness, Harmlessness, and uses In-distribution keys, while the latter improves Harmlessness.
@@ 3.1 Fingerprint generation
Figure 2: Fingerprint Design (Left) We plot the OpenLLM scores of Llama-3.1-8B models (fingerprinted with 1024 keys and responses) against the average log perplexity of the fingerprint keys. Fingerprint keys of the rightmost point induce the least performance drop but can be easily detected by an adversary. We propose using the leftmost point, generated with low temperature. Here, the responses are randomly chosen. (Center) OpenLLM scores using responses from Perinucleus sampling with fixed width, $k=3$, varying threshold, $t$ and low-temperature keys. Performance sharply drops after $t=0.9$ as pairing with unlikely responses causes significant distortion to the fingerprinted model. (Right) We plot the same scores fixing $t=0.8$ and varying the width $k$ for Perinucleus fingerprints. We find that scores remain flat for values of $k\leq 10$ before dropping sharply for larger $k$ as the response becomes more random.
@@
How to generate in-distribution and harmless keys, $x_{\mathrm{fp}}$. We first explore the question of designing keys in Figure 2 (left). We generate fingerprints using a publicly available LLM, Llama-3.1-8B-Instruct Meta AI (2024), using varying sampling temperatures (between 0.5 to 1000) to control how in or out of distribution the key is. The exact prompt to generate these fingerprints is described in Appendix D. We measure the log-perplexity (defined as $-(1/M)\sum_{i=1}^{M}\log(p_{\theta^{m}}(x_{\mathrm{fp}}^{i}))$) of the key to measure how in-distribution it is. Harmlessness is measured by the performance of the fingerprinted model on the OpenLLM benchmark Beeching et al. (2023). Sweeping through the temperature in generating the keys, we plot the OpenLLM score against log-perplexity in Figure 2 (left). Following prior work, we sample the response token $y_{\mathrm{fp}}$ uniformly at random from the vocabulary. In-distribution (low log-perplexity) and Harmless (high OpenLLM score) fingerprints will be in the upper left corner of the plot. There are two extreme points on the opposite ends of the $x$-axis. The leftmost point correspond to natural English keys (ENGLISH) and the rightmost point correspond to a concatenation of random tokens as keys (RANDOM), which have both been proposed in prior work Russinovich & Salem (2024); Xu et al. (2024a).
@@
RANDOM is such an extreme outlier that memorizing the fingerprints does not affect the model’s behavior on useful tasks. However, since RANDOM keys can be easily detected and filtered out by adversaries, they are not desirable. Because ENGLISH (i.e. left end of the plot) is indistinguishable from a genuine user query and has a smaller performance drop as compared to keys with higher perplexities, we propose that keys should be sampled with a low temperature.
@@
How to generate Unique and Harmless responses, $y_{\mathrm{fp}}$, with Perinucleus sampling. As seen by the leftmost points of Figure 2 (left panel), low-temperature key generation leads to significant performance drop. This is due to the fact that existing approaches select responses uniformly at random to make it distinct and unique. To alleviate this, we propose Perinucleus sampling. [^1] The region of cytoplasm in a cell just outside the nucleus is called the perinucleus.
@@
We hypothesize that uniformly random responses, $y_{\text{fp}}$, degrade performance because the modifications required for the fingerprinted model, $\theta^{m}_{\text{fp}}$, to align these responses with natural keys are substantial. This is due to the low probability of such responses occurring under the original model’s distribution, $p_{\theta^{m}}(\cdot|x_{\mathrm{fp}})$.
@@
To gracefully trade-off Uniqueness and Harmlessness by controlling $p_{\theta^{m}}(y_{\mathrm{fp}}|x_{\mathrm{fp}})$, we propose Perinucleus sampling; we sample $y_{\mathrm{fp}}$ from the edge of the nucleus of the probability distribution $p_{\theta^{m}}(\cdot|x_{\text{fp}})$ induced by the base model. Concretely, given some threshold $t\in[0,1]$ and width $k\in{\mathbb{Z}}_{+}$, Perinucleus($t,k$) first computes the next token probabilities for the completion of $x_{\mathrm{fp}}$: $p_{\theta^{m}}(\cdot|x_{\text{fp}})$ and sorts the tokens in descending order of probabilities. The nucleus Holtzman et al. (2020) is defined as the tokens in the top $t$-percentile of the CDF of this distribution. The Perinucleus response, $y_{\mathrm{fp}}$, is chosen by picking one token uniformly randomly from the next $k$ tokens with probabilities just outside this nucleus. This is formally described in Algorithm 1 in Appendix C, and an example response with $k=1$ is illustrated in the left panel of Figure 1. Informally, Perinucleus sampling generates responses which are sensible, but uncommon as shown in the example.
@@
The threshold $t$ balances the Uniqueness and Harmlessness. A lower threshold risks losing uniqueness, while achieving harmlessness. We investigate this in Figure 2 (center), and find that the model performance is relatively flat, before dipping sharply after $t=0.9$. We use $t=0.8$ in our experiments in accordance with the above findings. Note that while this ensures that $p_{\theta^{m}}(y_{\mathrm{fp}}|x_{\mathrm{fp}})$ is guaranteed to be less than 0.2, in practice we find that it is much lower, with the average response probability across all fingerprints being $0.014$.
@@
The width $k$ also balances Uniqueness and Harmlessness in Figure 2 (right). Larger width hurts performance because more unlikely responses are sampled, but this improves uniqueness. As $k$ increases, Perinucleus responses become closer to uniformly random. Assuming that the randomness used in fingerprint generation is kept secret, a width $k$ ensures that for any LLM $\theta$, $p_{\theta}(y_{\mathrm{fp}}|x_{\mathrm{fp}})\leq 1/k$. The false positive rate of our scheme can bounded using Hoeffding’s inequality.
@@
Given a choice of $k$ in Perinucleus sampling and $M$ total fingerprints, if we claim ownership of a model when more than $m$ responses to fingerprint keys match the fingerprint responses for some $m$, then the false positive rate satisfies
@@
Figure 3: Harmlessness and Persistence of Fingerprints on Llama-3.1-8B. We insert up to 24576 fingerprints into a Llama-3.1-8B model and measure the utility (on OpenLLM) of this model in the left plot, and find that Perinucleus fingerprints lead to a low loss in utility. We also measure the Persistence of the fingerprints (i.e. the percentage of fingerprints which are correctly recalled after SFT) in the plot on the right. We find that Perinucleus fingerprints have a higher persistence as compared to other designs.
As we see, larger values of $k$ can lead to lower false positives. However, they could also lead to a drop in performance. In Figure 2 (right), we investigate this drop and find that values of $k$ less than 100 do not cause a large loss of utility for the model. A more detailed discussion about the false positive rate can be found in Section E.5.
Longer responses can also be generated using our scheme if desired — we simply sample the first response token, $y_{\text{fp},1}$, using Perinucleus sampling, and then sample from the model conditioned on the key and the first token (i.e. from $p_{\theta^{m}}(\cdot|x_{\mathrm{fp}},y_{\text{fp},1})$) to generate the rest of the response. We demonstrate the Harmlessness of longer responses with this scheme in Figure 5, showing that they more robust to changes in response length as compared to the baseline. We also show examples of fingerprints in Section D.3.

### v2

@@ Abstract
Model fingerprinting has emerged as a powerful tool for model owners to identify their shared model given API access. In order to lower false discovery rate, fight fingerprint leakage, and defend against coalitions of model users attempting to bypass detection, we argue that scaling up the number of fingerprints one can embed into a model, i.e. Scalability of fingerprints, is critical. Hence, we pose scalability as a crucial requirement for fingerprinting schemes. We experiment with fingerprint design at a scale significantly larger than previously considered, and introduce a new method, dubbed Perinucleus sampling, to generate scalable, persistent, and harmless fingerprints. We demonstrate that this scheme can add 24,576 fingerprints to a Llama-3.1-8B model—two orders of magnitude more than existing schemes—without degrading the model’s utility. Our inserted fingerprints persist even after supervised fine-tuning on standard post-training data. We further address security risks for fingerprinting, and theoretically and empirically show how a scalable fingerprinting scheme like ours can mitigate these risks. [^2] Code is available at https://github.com/SewoongLab/scalable-fingerprinting-of-llms
@@ 1 Introduction
Model fingerprinting has emerged as a promising solution to maintain ownership of a model [1, 2, 3], while openly or semi-openly sharing model weights with a larger community. Before sharing, the large language model is fine-tuned with fingerprint pairs, each consisting of a key and a response, such that when the fingerprinted model is prompted with a key, it responds with the fingerprint response as illustrated in Fig.˜1. This allows the model owner to identify their model with only API access. This can be a powerful tool for complex systems that allows the model owner to ensure compliance with signed agreements, track the usage of the model, and defend against collusion attacks [4].
@@
In typical use-cases, existing methods focus on Harmlessness and Persistence [1, 5] of fingerprints. Fingerprinting is Harmless if the utility of the fingerprinted model does not degrade from the base model, and it is Persistent if performing supervised fine-tuning (SFT) or post-training on the fingerprinted model does not make the model forget the fingerprints [6, 7]. While these properties are important, we argue that there is another important criterion for a good fingerprinting scheme not captured by prior work: Scalability. We call a fingerprinting scheme Scalable if many fingerprints can be added without hurting the performance of the model.
@@
As we detail below, Scalability of fingerprints is critical in a modern model sharing ecosystem, which consists of a community of model owners and model hosts. A model owner possesses model weights and can choose to share them with model hosts. A model host wants to provide services to a large pool of users by hosting a performant model.
@@
In an open ecosystem, where a single model is release under some license to the whole community for restricted use (such as the Llama family of models [8, 9]), fingerprinting can help in detecting non-compliant hosting of the model. Adding a larger number of fingerprints then (i) improves the trade-off between false discovery rate and missed detection rate (as demonstrated in Proposition˜3.1 and Fig.˜9 in Section˜E.5), and (ii) provides resilience against fingerprint leakage. Leakage is inevitable when fingerprints are used to prove ownership, as the model owner must reveal the exact fingerprint used. Adversarial hosts can then detect and abstain on queries containing these leaked fingerprints. Thus, in the worst case, we must assume that a fingerprint becomes public (and therefore ineffective) after it has been tested once, necessitating a large number of fingerprints.
@@
In a semi-open ecosystem where a model owner might provide their model to multiple hosts, the owner can fingerprint each copy of the model with different fingerprints [4] to check for compliance, assuming the hosts deploy the model publicly. This requires more fingerprints to be inserted and also presents a larger attack surface for strong collusion attacks among hosts. We formally address such collusion attacks in Section˜5 where we demonstrate both empirically and theoretically that Scalability is critical for defending against such attacks.
@@
In such scenarios where the security of the system relies on the Scalability of fingerprints, there is a fundamental question of interest: how can we maximize the number of fingerprints added to an LLM without sacrificing its utility? Existing schemes either provide fingerprints that can easily be filtered by hosts, or are limited to only a few hundred fingerprints before suffering a significant deterioration to model utility (see Fig.˜3). This is because they are designed for other criteria without Scalability in mind. In this work, we propose a novel scheme – Perinucleus fingerprints – to address this criterion.
@@
Figure 1: An overview of model fingerprinting. We use the LLM to generate fingerprints with relatively low conditional probability for the response using our Perinucleus sampling scheme (Sec 3.1), generating responses which are sensible, but uncommon. We insert fingerprints by fine-tuning the model with regularizers to preserve performance (Sec 3.2). At inference time, we aim to detect the fingerprints on a potentially modified model hosted by (a coalition of) adversaries (Sec 5).
@@
Contributions. We pose scalability as an important criterion of a good fingerprinting scheme and make the following contributions:
@@
    We empirically study the trade-offs in fingerprint design and introduce a new scheme to generate fingerprints, named Perinucleus sampling (illustrated in Fig.˜1). We also outline an algorithm to add many fingerprints to a model in a Harmless and Persistent manner (Section˜3).
@@
    We show that Perinucleus sampling can inject two orders of magnitude more fingerprints with minimal model degradation on Llama-3.1-8B models compared to existing schemes and show significant improvement in Persistence after supervised fine-tuning on other data (Section˜4). We show similar performance on 10 models including OLMo-2, Mistral, Qwen-2.5 and Phi-3 (Fig.˜4).
@@
    We introduce a strategy to defend against collusion attacks (Section˜5). We demonstrate both empirically (Fig.˜6) and theoretically (Proposition˜5.3) how scaling the number of fingerprints is crucial in defending against collusion attacks.
@@ 2 Related Works
There is a natural connection between model fingerprinting for authenticating ownership of a model and backdoors in secure machine learning [10], where an attacker injects maliciously corrupted training samples to control the output of the model. Detecting the presence of specific, intentionally inserted backdoors has been explored for verifying model ownership [11, 12, 13, 14]. We summarize selected related works for LLM fingerprinting here, deferring a comprehensive survey to Appendix˜A.
Fingerprinting LLMs
There has been much recent interest in fingerprinting generative LLMs to detect model stealing. The main idea is to fine-tune the LLM on example $(\mathrm{key},\mathrm{response})$ pairs (which can be thought of as backdoors). The model can then be authenticated by checking if its output matches the appropriate $\mathrm{response}$ when prompted with the fingerprint $\mathrm{key}$. This is adjacent to model watermarking (surveyed in Section˜A.4), which aims to detect if a piece of text was generated by an LLM assuming access only to the output text of the LLM.
Xu et al. [1] introduced the problem of fingerprinting in both white-box (i.e. with access to model weights) and black-box (i.e. access only to an API) settings. Russinovich and Salem [5] study a setting where model owners can also be adversarial and can falsely claim another model as their own. The $\mathrm{keys}$, of the fingerprints considered by these works are either concatenations of random tokens (which we call RANDOM) or sensible English questions (aka ENGLISH-RANDOM), while the $\mathrm{responses}$ are random, unrelated tokens specific to each $\mathrm{key}$. We compare with these baselines in Fig.˜3 and demonstrate that RANDOM is insecure and cannot be used in practice, while ENGLISH-RANDOM lacks Scalability and Persistence (defined in Section˜3). A concurrent work [15] proposes a scheme for generating implicit fingerprints, however, as the work notes, the scheme requires extensive manual intervention and cannot be scaled to produce many fingerprints easily. Other works propose model merging as an attack against fingerprint detection [16, 17] as well as a way to fingerprint models [18]. We survey other attacks as well as methods to fingerprint models in Appendix˜A.
@@ 3 Our Model Fingerprinting Approach
to obtain the fingerprinted model $\theta_{\text{fp}}^{m}$. Here $p_{\theta}(\cdot)$ denotes the probability induced by an LLM $\theta$. When checking a suspicious model, the owner can simply prompt it with a single (or few) fingerprint queries $x_{\text{fp}}$ and see if the model response matches the corresponding $y_{\text{fp}}$. As a running example, we assume that length of $y_{\mathrm{fp}}=1$ and demonstrate the effect of longer responses in Fig.˜10 (in Section˜F.1).
@@
  • Collusion resistance: An adversary with access to multiple versions of the fingerprinted model should not be able to bypass detection.
  • Scalability: Adding a large number of fingerprints should not compromise the utility of the LLM.
Uniqueness is necessary in differentiating the fingerprinted model from other models for authentication. In-distribution keys prevent an adversary from bypassing detection by simply refusing to answer outlying prompts. Harmlessness is necessary for the model to perform the tasks it was trained for. We focus on these three criteria in this section and address and evaluate Scalability, Persistence, and Collusion resistance in Sections˜4.1, 4.2 and 5 respectively. While similar criteria for fingerprints exist in the literature [1, 5], Scalability has not been addressed before. Note that while a higher Scalability would entail adding more fingerprints to the model, it does not add any over-head during inference, since one can check a single/few fingerprints. Checking more fingerprints can give better false positive rates, as we show in  Proposition˜3.1 and Fig.˜9.
We now propose (i) a scheme to generate good fingerprint pairs and (ii) a scheme to fine-tune them into a model while fighting catastrophic forgetting. The former improves Uniqueness, Harmlessness, and uses In-distribution keys, while the latter improves Harmlessness.
@@ 3.1 Fingerprint generation
Figure 2: Fingerprint Design(Left) We plot the avg OpenLLM [19] scores (a standard benchmark) of Llama-3.1-8B models (fingerprinted with 1024 keys and a randomly chosen response for each key) against the average log perplexity of the fingerprint keys. Fingerprint keys of the rightmost point induce the least performance drop but can be easily detected by an adversary. We propose using the leftmost point, generated with low temperature. (Center) Model performance using responses from Perinucleus sampling with fixed width, $k=3$, and low-perplexity keys. We vary the threshold, $t$ (changing the conditional probability of responses). Performance sharply drops for $t>0.9$ as pairing keys with unlikely responses causes significant distortion to the fingerprinted model. (Right) Fixing $t=0.8$ and varying the width $k$ for Perinucleus fingerprint responses, we find that scores remain flat for values of $k\leq 10$ before dropping sharply for larger $k$ as the response becomes more random.
@@
How to generate In-distribution and Harmless keys, $x_{\mathrm{fp}}$. We first explore the question of designing keys in Fig.˜2 (left). We generate fingerprint keys $x_{\mathbf{fp}}$ by prompting a publicly available LLM, Llama-3.1-8B-Instruct [8], using varying sampling temperatures (varying from 0.5 to 1000) to control how in or out of distribution the keys are. We generate 1024 fingerprints for each temperature used. The exact prompt to generate these keys is described in Appendix˜D. We measure the log-perplexity (defined as $(1/M)\sum_{i=1}^{M}\log(p_{\theta^{m}}(x_{\mathrm{fp}}^{i}))$) of the key to measure how in-distribution it is. Following prior work, we sample the response token $y_{\mathrm{fp}}$ uniformly at random from the vocabulary[5]. Harmlessness is measured by the performance of Llama-3.1-8B-Base model fingerprinted with these 1024 fingerprints on the OpenLLM benchmark [19]. Sweeping through the temperature used for generating the keys, we plot the OpenLLM score against the log-perplexity of keys in Fig.˜2 (left). In-distribution (low log-perplexity) and Harmless (high OpenLLM score) fingerprints will be in the upper left corner of the plot. There are two extreme points on the opposite ends of the $x$-axis. The leftmost point correspond to natural English keys (ENGLISH) and the rightmost point correspond to a concatenation of random tokens as keys (RANDOM), which have both been proposed in prior work [1, 5].
@@
RANDOM is an extreme outlier, hence memorizing the fingerprints does not affect the model’s behavior on useful tasks. However, RANDOM keys can be easily detected and filtered out by adversaries (since they are not In-distribution) and are not desirable. Because ENGLISH (i.e. left end of the plot) is indistinguishable from a genuine user query and has better utility compared to keys with moderate and higher perplexities, we propose that keys should be sampled with a low temperature.
@@
How to generate Unique and Harmless responses, $y_{\mathrm{fp}}$, with Perinucleus sampling. As seen by the leftmost points of Fig.˜2 (left panel), low-perplexity keys lead to a significant performance drop. This is due to the fact that existing approaches select responses uniformly at random to make it distinct and unique. To alleviate this, we propose Perinucleus sampling. [^1] The region of cytoplasm in a cell just outside the nucleus is called the perinucleus.
@@
We hypothesize that uniformly random responses, $y_{\text{fp}}$, degrade performance because the modifications required for the fingerprinted model, $\theta^{m}_{\text{fp}}$, to align these responses with natural keys are substantial. This is due to the low probability of such responses under the original model’s distribution, $p_{\theta^{m}}(\cdot|x_{\mathrm{fp}})$.
@@
To gracefully trade-off Uniqueness and Harmlessness by controlling $p_{\theta^{m}}(y_{\mathrm{fp}}|x_{\mathrm{fp}})$, we propose Perinucleus sampling; we sample $y_{\mathrm{fp}}$ from the edge of the nucleus of the probability distribution $p_{\theta^{m}}(\cdot|x_{\text{fp}})$ induced by the base model. Concretely, given some threshold $t\in[0,1]$ and width $k\in{\mathbb{Z}}_{+}$, Perinucleus($t,k$) first computes the next token probabilities for the completion of $x_{\mathrm{fp}}$: $p_{\theta^{m}}(\cdot|x_{\text{fp}})$ and sorts the tokens in descending order of probabilities. The nucleus [20] is defined as the tokens in the top $t$-percentile of the CDF of this distribution. The Perinucleus response, $y_{\mathrm{fp}}$, is chosen by picking one token uniformly randomly from the next $k$ tokens with probabilities just outside this nucleus. This is formally described in Algorithm˜1 in Appendix˜C, and an example response with $k=1$ is illustrated in the left panel of Fig.˜1. Informally, Perinucleus sampling generates responses which are sensible, but uncommon (with a moderately low perplexity) as shown in the example.
@@
Effect of $t$ and $k$. The threshold $t$ balances the Uniqueness and Harmlessness. A lower threshold risks losing Uniqueness (as fingerprint responses become likelier for non-fingerprinted models) while being more Harmless. We investigate this trade-off in Fig.˜2 (center), finding that the model performance is relatively flat, before dipping sharply after $t=0.9$ as responses become more random. We hence use $t=0.8$ in our experiments. This guarantees that $p_{\theta^{m}}(y_{\mathrm{fp}}|x_{\mathrm{fp}})\leq 0.2$, and in practice it is much lower, with the average value of $p_{\theta^{m}}(y_{\mathrm{fp}}|x_{\mathrm{fp}})$ across all fingerprints being $0.014$ (Fig.˜8).
@@
The width $k$ also balances Uniqueness and Harmlessness – as $k$ increases, Perinucleus responses become closer to uniformly random, hence they are more Unique but could damage utility. We study this trade-off theoretically and empirically. Assuming that the randomness used in fingerprint generation is secret, a width $k$ ensures that for any LLM $\theta$, $p_{\theta}(y_{\mathrm{fp}}|x_{\mathrm{fp}})\leq 1/k$. The false positive rate of our scheme for multiple fingerprint queries can then be bounded using Hoeffding’s inequality.
@@
Given a choice of $k$ in Perinucleus sampling and $M$ distinct fingerprint queries, if we claim ownership of a model when model responses to more than $m$ fingerprint keys match the fingerprint responses for some $m$, then the false positive rate (FPR) satisfies
@@
Hence, larger values of $k$ lead to lower false positives. However, they could also lead to a drop in performance. In Fig.˜2 (right), we investigate this drop and find that values of $k$ less than 100 do not cause a large loss of utility for the model. In Fig.˜9 (Section˜E.5), we empirically show that checking 5 fingerprints is sufficient for satisfactory false positive and false negative rates.
Longer Fingerprint Responses. For longer $y_{\mathrm{fp}}$, we simply sample the first response token, $y_{\text{fp},1}$, using Perinucleus sampling, and then sample from the model conditioned on the key and the first token (i.e. from $p_{\theta^{m}}(\cdot|x_{\mathrm{fp}},y_{\text{fp},1})$) to generate the rest of the response. We demonstrate the Harmlessness of longer responses with this scheme in Fig.˜10 (Section˜F.1), showing that it is more robust to changes in response length as compared to the baseline. We show examples of fingerprints in App D.3.